OpenStudy (diyadiya):

Find the intervals in which the function f given by f(x)=sinx + cosx ,o≤x≤2pi is strictly increasing or strictly decreasing

OpenStudy (diyadiya):

\[f(x)=sinx+cosx\]\[f'(x)=cosx-sinx\]

OpenStudy (mr.math):

Can you solve \(\cos(x)-\sin(x)=0\), as we just did?

OpenStudy (anonymous):

sin(x) = -cos(x) sin(x)/-cos(x) = 1 tan(x) = -1 which gives you the solution 3pi/4, 5pi/4

OpenStudy (anonymous):

Take the inequalities at these two points

OpenStudy (diyadiya):

\[Cos(x)=Sin(x) \] \[x= \frac{\pi}{4}, \frac{5\pi}{4}\]

myininaya (myininaya):

f'(x)=cos(x)-sin(x) set =0 cos(x)-sin(x)=0 => cos(x)=sin(x) this happens at x=pi/4, pi/4+pi=pi/4,5pi/4 | ------|-------|-------| 0 pi/4 5pi/4 2pi choose test points in each of these intervals to plug into f' to see if you get negative output or positive output a positive output=> it is increasing on that interval a negative output=> it is decreasing on that interval

myininaya (myininaya):

gj diyadiya

OpenStudy (diyadiya):

Thanks :) Now?

OpenStudy (anonymous):

Nice job!

OpenStudy (mr.math):

Amazing! Well, myin had everything explained, there's nothing more to say! :)

OpenStudy (anonymous):

so you should get 3pi/4 + 2pi(k) and 5pi/4 + 2pi(k) K being an element of all intergers so you would set it up like this o≤ (3pi/4) +2pi(k) ≤ 2pi you should be left with k and two numbers it fits inside figure out what integers are present in the inequality then sub them into the original domain (3pi/4) +2pi(k) Do with the other one and you will find all possible solutions

OpenStudy (diyadiya):

OK letme go through it!

OpenStudy (turingtest):

@ cuddlypony it will be pi/4 and 5pi/4 we are solving sinx=cosx this time

OpenStudy (anonymous):

oh sorry I miss read

OpenStudy (diyadiya):

Can you show the last part ?

OpenStudy (anonymous):

no it wouldn't turing test sin(x) + cos(x) = f(x) or did I mess up

OpenStudy (anonymous):

So what integers did you get for k

myininaya (myininaya):

we want to find where f is decreasing and increasing what function will tell us where f is increasing and decreasing@ cuddle?

OpenStudy (diyadiya):

\[(0,\pi/4 ) (\pi/4,5\pi/4)(5\pi/4,2\pi)\]

myininaya (myininaya):

f tells us where f' tells us the increasing/decreasing part f'' tells us about concavity

myininaya (myininaya):

we need f' here so f'(x)=cos(x)-sin(x)

OpenStudy (anonymous):

sorry i read the question wrong disregard my input :l

myininaya (myininaya):

set this puppy =0

OpenStudy (turingtest):

good Diya, so pick a point in each interval and see if f'(x) is >0 or <0

OpenStudy (mr.math):

Okay! In general a function is decreasing when its derivative is negative and it's increasing when the derivative is positive. In our case, we have found three intervals of signs for the derivative and we need to test each of them. Following so far?

OpenStudy (diyadiya):

Yup!

OpenStudy (diyadiya):

So i should plug in the values of each interval to "cos(x)-sin(x) right ?

OpenStudy (mr.math):

It's very easy to test the sign of each given interval: choose any value for x inside the interval (not in the boundaries) and substitute into f'(x) and see what you get.

OpenStudy (mr.math):

Yes Diya.

OpenStudy (diyadiya):

Ok 1 min

OpenStudy (diyadiya):

cos(pi/6)-sin(pi/6) = 1/2 -sqrt3/2 is it like this ? o.O

OpenStudy (mr.math):

I think you switched the values :D

OpenStudy (diyadiya):

oh yeah

OpenStudy (diyadiya):

sqrt3/2-1/2

OpenStudy (turingtest):

now is that >0 or <0 ???

OpenStudy (diyadiya):

>0

OpenStudy (mr.math):

\(\cos{\frac{\pi}{6}}-\sin{\frac{\pi}{6}}=\frac{\sqrt{3}}{2}-\frac{1}{2}>0\), since \(\sqrt{3}>1\).

OpenStudy (mr.math):

So, you conclude that..?

OpenStudy (diyadiya):

Strictly Increasing

OpenStudy (mr.math):

on..?

OpenStudy (diyadiya):

(0,pii/4)

OpenStudy (mr.math):

Good! when you say Strictly use open intervals.

OpenStudy (diyadiya):

okay 1min brb sorryy

OpenStudy (mr.math):

Alright! Apply the same idea to the other two intervals.

OpenStudy (diyadiya):

ThankYouu Myininaya ,Mr.Math & turingtest :D

OpenStudy (mr.math):

Glad to help!

OpenStudy (turingtest):

anytime :)